mhchen:

Find variance of uniform distribution on [0,1] using moment generating function

mhchen:

\[M(t) = E(e^{tx}) = \int\limits_{0}^{1}e^{tx}dx = \frac{e^{t}-1}{t}\] I checked this and it's correct. Now then: \[M'(t) = \frac{te^t-e^t+1}{t^2}\] so \[E(X) = M'(0) = \lim_{t \rightarrow 0}\frac{te^t-e^t+1}{t^2} = \frac{1}{2}\] Variance formula is \(E(X^2) - E(X)^2\) And \[E(X^2) = M''(0)\] and we calculate \[M''(t) = \frac{t^2e^t-2te^t+2e^t-2}{t^3}\] and \[M''(0) = \lim_{t \rightarrow 0}\frac{t^2e^t-2te^t+2e^t-2}{t^3} = 0\] But then \[E(X^2) - E(X)^2 = 0 - \frac{1}{4} = -\frac{1}{4}\] which can't be right. I can show my work for every step but it takes too long to type so if there's one part where you think it's wrong please tell me.

justjm:

Hopefully I can check your computational work. Let me take a look

mhchen:

BRO I AM AN IDIOT I calculated M''(0) wrong

mhchen:

M''(0) is 1/3 not 0

mhchen:

basically I was trying to use l'hospital's rule but instead of doing the derivative of top and bottom, I did the derivative of the entire thing instead

justjm:

Unless i'm wrong why did you take the limit of M'(t) approaching 0 to find M'(0) if you had M'(t)? Other than that the lhopitals seems good

justjm:

Oh lmao I was going to ask

mhchen:

cuz it's undefined at 0

justjm:

Oh yeah got it. didn't see there

justjm:

Okay..I got the same for the second derivative. let me see the M''(0)

justjm:

I got 1/3 ur right

justjm:

Is it supposed to be 1/3? I don't know the stat part, I don't think AP stat is enough for this

mhchen:

Nah AP stats is not enough. You also need calculus

mhchen:

As long as you learn how to use this symbol: \[\int\limits_{a}^{b}f(x)dx\] then that's enough

justjm:

definitely bruv